A thermostated chamber at a height h above the earth surface maintained at 30 degree has a clock fitted with uncompensated pendulum. The maker of the clock for chamber mistakenly designed it to maintain correct time at 20degree. It is found that when the chamber is brought to the earth surface the clock in it clicked correct time. Re is the radius of earth. The coefficient of the material of the pendulum is A) h/Re b) h/5Re c) 5Re/h d) Re/h

Grade:12

1 Answers

Vikas TU
4 years ago
Variaton of g with altitude = g [1/2h/R]
delta g = 2hg/R
Linear expansively = Delta l / l*t
T= 2*pie*sqrt(l/g)
Loss of time due to change in length = gain in time due to change in g 
1 + 0.5 delta l /l = delta g / g 
The cofficient of material of pendulum = h/5Re